OpenStudy (anonymous):

Find dy/dx. y=9^(-x)

OpenStudy (kira_yamato):

dy/dx = 9^x ln 9 (-1) \[\frac{dy}{dx} = -2\left( 9^x \right)\ln3\]

OpenStudy (anonymous):

You mean 9^(-x)?

OpenStudy (kira_yamato):

yeah... gomen...

OpenStudy (anonymous):

So does du/dx = -x^(-2)?

OpenStudy (kira_yamato):

What is u?

OpenStudy (anonymous):

Chain rule?

OpenStudy (anonymous):